Please help with the following problem?
Question: solve the system of equations algebraically?
#12. -5x+2y = 8
y = -3x+26

Answer by uma(370)

The given equations are:
-5x + 2y = 8
y = -3x + 26
Rearranging (adding 3x to both the sides of the above equation) we get..
3x + y = 26-------------(1)
-5x + 2y = 8---------------(2)
I choose to elimonate y.
(1) multiplied by 2 gives.. 6x + 2y = 52 ------------(3)
-5x + 2y = 8 ------------(2)
eqn(3) - eqn (2) gives us..
6x + 2y -(-5x + 2y) = 52 - 8
==> 6x + 2y + 5x - 2y = 44 [Removing the parenthesis]
==> 11x = 44 [Adding the like terms]
==> x = 4 [dividing by 11 on both the sides]
Substituting x = 4 in eqn(1) we get..
3(4) + y = 26
==> 12 + y = 26
==> y = 26 - 12 [subtracting 12 from both the sides of the equation]
==> y = 14
THus (x, y) = (4, 14)
Good Luck!!!